How do I write the product/quotient in standard form for these problems 1+3i/2i and 2i/3+i?
Answer by Alan3354(30983)

How do I write the product/quotient in standard form for these problems 1+3i/2i and 2i/3+i?
============
Assuming it's (1+3i)/2i
Multiply NUM and DEN by i
= i(1+3i)/2i^2
= i(1+3i)/(-2)
= -(i+3i^2)/2
= -(i-3)/2
= (3-i)/2
====================
2i/(3+i) Use parentheses for clarification.
Multiply NUM and DEN by the conjugate of the DEN (3-i)
Try it.
email via the Thank You note if you want me to check your work.